# Enhancing Layer Attention Efficiency through Pruning Redundant Retrievals

### v3

@@ 4.1 Image Classification
We carried out experiments on the CIFAR-10, CIFAR-100, and ImageNet-1K Deng et al. (2009) datasets, utilizing ResNet He et al. (2016) as the backbone compared with various baselines, such as channel attention methods, spatial attention models, and other layer interaction methods. For the CIFAR-10 and CIFAR-100 datasets, we applied the data augmentation techniques described in Huang et al. (2016). For the ImageNet-1K dataset, the augmentation techniques and hyperparameter configurations followed those specified by He et al. (2016, 2019).
@@
Results on CIFAR-10 and CIFAR-100 ELANet, which integrates the proposed method of pruning attention operations from certain layers, demonstrates a clear advantage over baseline models on both CIFAR-10 and CIFAR-100. As shown in Table 1, ELA achieves 92.45% and 71.67% top-1 accuracy on the ResNet-20 backbone, surpassing ResNet-20 by 1.10% and 3.72% on CIFAR-10 and CIFAR-100, respectively. ELA also outperforms several channel and spatial attention models, such as SE Hu et al. (2018), BAM Park (2018), and CBAM Woo et al. (2018). In the ResNet-56 backbone, ELA surpasses SE, BAM, and CBAM by 0.59%, 0.71%, and 0.39% on CIFAR-10, and by 0.53%, 0.67%, and 0.60% on CIFAR-100, respectively, with lower parameters introduced. Furthermore, in the ResNet-110, our model exceeds a number of layer interaction models, surpassing DIAHuang et al. (2020) and MRLA-L Fang et al. (2023) by 0.39% and 0.38% on CIFAR-10, and by 0.21% and 0.33% on CIFAR-100, respectively. These results highlight ELA’s superior classification performance while maintaining efficient parameter usage across various network depths.
@@
ELA significantly reduces training time. As shown in Figure 4, the training time for the three networks on CIFAR-10 and CIFAR-100 are nearly identical for one epoch, with MRLA-B exhibiting the longest training time. Compared to MRLA-B, MRLA-L reduces training time by approximately 10%-15%, while the proposed ELA achieves an even greater reduction, decreasing training time by around 30%-35% relative to the original MRLA-B. Moreover, the reduction in training time becomes more significant as the number of network layers increases.
Data | CIFAR-10 | CIFAR-100 | ||
Models | Params | Top-1 | Params | Top-1 |
ResNet-20 | 0.22M | 91.35 | 0.24M | 67.95 |
SE | 0.24M | 91.22 | 0.27M | 68.61 |
ECA | 0.22M | 91.45 | 0.24M | 67.35 |
DIA | 0.44M | 92.02 | 0.46M | 68.83 |
CBAM | 0.24M | 91.97 | 0.27M | 67.71 |
BAM | 0.27M | 91.87 | 0.29M | 67.91 |
MRLA-L | 0.23M | 92.21 | 0.25M | 70.84 |
ELA | 0.23M | 92.45 | 0.25M | 71.67 |
ResNet-56 | 0.59M | 93.01 | 0.61M | 72.36 |
SE | 0.66M | 93.56 | 0.68M | 73.78 |
ECA | 0.59M | 94.11 | 0.61M | 73.38 |
DIA | 0.81M | 93.56 | 0.83M | 74.01 |
CBAM | 0.66M | 93.76 | 0.68M | 73.71 |
BAM | 0.73M | 93.44 | 0.75M | 73.64 |
MRLA-L | 0.62M | 93.88 | 0.64M | 74.13 |
ELA | 0.62M | 94.15 | 0.64M | 74.31 |
ResNet-110 | 1.15M | 92.94 | 1.17M | 73.04 |
SE | 1.28M | 94.30 | 1.30M | 75.18 |
ECA | 1.15M | 94.27 | 1.17M | 75.08 |
DIA | 1.37M | 94.13 | 1.39M | 75.14 |
CBAM | 1.28M | 93.79 | 1.30M | 74.87 |
BAM | 1.42M | 93.84 | 1.45M | 74.99 |
MRLA-L | 1.21M | 94.14 | 1.23M | 75.02 |
ELA | 1.21M | 94.52 | 1.23M | 75.35 |
Results on ImageNet-1K We compare ELA with a range of state-of-the-art models, using ResNet as backbone. As shown in Table 2, ELA achieves 77.9% Top-1 accuracy on the ResNet-50, an improvement of 1.8% over its backbone. Compared to attention-based models like SE and CBAM, ELA offers 1.2% and 0.6% higher Top-1 accuracy, respectively, while using a smaller number of parameters. Moreover, ELA outperforms Chen et al. (2018) by 0.9%, AA Bello et al. (2019) by 0.2%, and GC Cao et al. (2019) by 0.2%. Furthermore, ELA performs better than MRLA-L, which achieves 77.7%, by 0.2% in the accuracy of Top-1. For the deeper ResNet-101 backbone, ELA again delivers the best performance, achieving 78.8% in Top-1 accuracy, surpassing ResNet-101 by 1.4%. It also outperforms SE, CBAM, and AA by 0.2%, 0.3%, and 0.1%, respectively. Notably, ELA achieves 0.1% higher Top-1 accuracy than MRLA-L, which reaches 78.7%. When compared to lightweight models such as ECA, ELA provides an additional 0.4% improvement in Top-1 accuracy for ResNet-50 and 0.1% for ResNet-101, despite a slight increase in parameters. The results demonstrate that ELA significantly outperforms other models in terms of Top-1 and Top-5 accuracy, while maintaining competitive computational efficiency.
A significant reduction in training time for ELA on the ImageNet-1K dataset is presented in Table 3. ELA reduces the training time per epoch by 29.7% compared to MRLA-B for ResNet-50 and by 32.6% for ResNet-101, outperforming MRLA-L, which achieves reductions of only 13.5% and 16.3% for ResNet-50 and ResNet-101, respectively. Furthermore, as the number of network layers increases, the proportion of training time reduction compared with MRLA-B becomes more pronounced.
@@
ELA | 44.9M | 8.1B | 78.8 | 94.6 |
[…]
MRLA-L | 3.2 | |
ELA | 2.6 | |
[…]
MRLA-L | 7.2 | |
[…]
4.2 Object Detection
Experimental Settings In the object detection task, our approach was evaluated on the COCO2017 dataset using Faster R-CNN and Mask R-CNN as detectors. All models were implemented using the open-source MMDetection toolkit Chen et al. (2019) to ensure consistency and reproducibility, with all configurations following the default MMDetection settings. Each detector underwent 12 training epochs, with KL divergence measured at the 1st and 7th epochs to analyze layer significance at early and later training stages. The hyperparameters and were set to 2 and 5, respectively, while was fixed at 0.2.
[…]
Fixed Layer Pruning | 73.790.30 | |
EBQM | 74.310.24 | |
[…]
Fixed Layer Pruning | 74.920.41 | |
[…]
@@ 4.2 Object Detection
Results on COCO2017 Our proposed ELA method achieves significant improvements in object detection performance compared to several state-of-the-art methods across various backbone networks and detectors. As highlighted in Table 5, when Faster R-CNN is used with ResNet-50, ELA outperforms SE, ECA, and RLA Zhao et al. (2021) by achieving 40.5% in AP, a gain of 2.8% over SE, 2.5% over ECA and 1.7% over RLA. It also shows improvements in AP50 and AP75, highlighting its superiority in detecting objects across different scales. With ResNet-101, ELA continues to show competitive performance, surpassing not only MRLA-L, but also SE, ECA, and other models, with a 0.1% improvement in AP and a notable increase in AP50 and AP75. Similarly, when Mask R-CNN is used as the detector, ELA maintains its edge, achieving a 0.2% increase in AP compared to MRLA-L on ResNet-50, and showing substantial improvements over methods like NL Wang et al. (2018) and GC Cao et al. (2019). These results confirm that ELA consistently enhances object detection accuracy, offering a balanced trade-off between performance and computational efficiency.
@@ 5.1 Evaluation of Different Layer Pruning Methods
To evaluate the effectiveness of EBQM, we conducted comparative experiments with two alternative strategies. First, a fixed layer pruning strategy was employed, where the number of pruned layers was determined by a hyperparameter. The pruned layer count was aligned as closely as possible with the number determined by Algorithm 1. Second, a threshold-based pruning strategy was implemented, using a fixed hyperparameter to define a KL divergence threshold: layers with values below the threshold were discarded, while those above were retained. The ablation experiments were conducted on CIFAR-100 using ResNet He et al. (2016) architectures of varying depths as baselines. Each experiment was repeated five times, and results were reported as Accuracy Standard Deviation to evaluate the model’s robustness and performance.
As shown in Table 4, it is evident that as the network depth increases, the instability of the Threshold-Based method also increases. The standard deviation rises from 0.44% in ResNet-56 to 0.79% in ResNet-110. A similar trend is observed with the Fixed Layer Pruning method. In contrast, the Top-1 accuracy achieved by our EBQM is consistently higher than that of the other two methods, reaching 74.31% and 75.35% for ResNet-56 and ResNet-110, respectively.
@@ 5.2 Evaluation of Different Distributions and Activation Functions in EBQM
The Beta distribution outperforms both the Normal and Gamma distributions, as well as the Softmax and Sigmoid functions, in terms of Top-1 accuracy, achieving 74.31% and 75.35% for ResNet-56 and ResNet-110, respectively, as shown in Table 6. Furthermore, the Beta distribution exhibits superior stability, with the lowest standard deviations of 0.24% and 0.22%, indicating more consistent performance compared to the other distributions.
@@ Appendix A Quantile Mapping
Algorithm 2 outlines the original Quantile Mapping Algorithm as introduced by Qian and Chang [2021]. In this method, the Gamma distribution and Exponential distribution are frequently utilized to model the target distribution . Their cumulative distribution functions (CDFs) are as follows:
@@
To assess the performance of EBQM in comparison to the existing Quantile Mapping Algorithms, we compared EBQM with the widely used Gamma Quantile Mapping (GQM) Piani et al. [2010] and Exponential Quantile Mapping (EQM) methods Heo et al. [2019].
[…]
EQM | 73.930.34 | |
EBQM | 74.310.24 | |
[…]
EQM | 75.050.37 | |
[…]
EBQM surpasses both GQM and EQM in Top-1 accuracy, achieving 74.31% and 75.35% for ResNet-56 and ResNet-110, respectively, as demonstrated in Table 7. These results surpass those of GQM, with accuracies of 74.02% and 74.98%, and EQM, with accuracies of 73.93% and 75.05%. Moreover, EBQM demonstrates the lowest standard deviation, at 0.24% and 0.22%, indicating superior stability when compared to GQM and EQM.

### v4

@@ 4.1 Image Classification
We carried out experiments on the CIFAR-10, CIFAR-100, and ImageNet-1K [6] datasets, utilizing ResNet [10] as the backbone compared with various baselines, such as channel attention methods, spatial attention models, and other layer interaction methods. For the CIFAR-10 and CIFAR-100 datasets, we applied the data augmentation techniques described in [14]. For the ImageNet-1K dataset, the augmentation techniques and hyperparameter configurations followed those specified by [10, 11].
@@
| Data | CIFAR-10 | CIFAR-100 | |||
| Models | Params | Top-1 | Params | Top-1 | |
| ResNet-20 | 0.22M | 91.35 | 0.24M | 67.95 | |
| SE | 0.24M | 91.22 | 0.27M | 68.61 | |
| ECA | 0.22M | 91.45 | 0.24M | 67.35 | |
| DIA | 0.44M | 92.02 | 0.46M | 68.83 | |
| CBAM | 0.24M | 91.97 | 0.27M | 67.71 | |
| BAM | 0.27M | 91.87 | 0.29M | 67.91 | |
| MRLA-L | 0.23M | 92.21 | 0.25M | 70.84 | |
| CAA | 0.59M | 92.32 | 0.61M | 71.21 | |
| SHSA | 0.41M | 92.27 | 0.43M | 71.64 | |
| DLA | 0.41M | 92.68 | 0.43M | 71.44 | |
| ELA | 0.23M | 93.13 | 0.25M | 72.07 | |
| ResNet-56 | 0.59M | 93.01 | 0.61M | 72.36 | |
| SE | 0.66M | 93.56 | 0.68M | 73.78 | |
| ECA | 0.59M | 94.11 | 0.61M | 73.38 | |
| DIA | 0.81M | 93.56 | 0.83M | 74.01 | |
| CBAM | 0.66M | 93.76 | 0.68M | 73.71 | |
| BAM | 0.73M | 93.44 | 0.75M | 73.64 | |
| MRLA-L | 0.62M | 93.88 | 0.64M | 74.13 | |
| CAA | 1.70M | 94.17 | 1.72M | 74.38 | |
| SHSA | 1.15M | 94.09 | 1.17M | 74.23 | |
| DLA | 0.80M | 94.52 | 0.82M | 74.79 | |
| ELA | 0.62M | 94.89 | 0.64M | 75.48 | |
| ResNet-110 | 1.15M | 92.94 | 1.17M | 73.04 | |
| SE | 1.28M | 94.30 | 1.30M | 75.18 | |
| ECA | 1.15M | 94.27 | 1.17M | 75.08 | |
| DIA | 1.37M | 94.13 | 1.39M | 75.14 | |
| CBAM | 1.28M | 93.79 | 1.30M | 74.87 | |
| BAM | 1.42M | 93.84 | 1.45M | 74.99 | |
| MRLA-L | 1.21M | 94.14 | 1.23M | 75.02 | |
| CAA | 3.37M | 94.22 | 3.39M | 75.22 | |
| SHSA | 2.27M | 93.76 | 2.29M | 75.31 | |
| DLA | 1.39M | 94.36 | 1.41M | 75.69 | |
| ELA | 1.21M | 95.08 | 1.23M | 76.25 | |
@@
Results on CIFAR-10 and CIFAR-100 ELANet, which integrates the proposed method of pruning attention operations from certain layers, demonstrates a clear advantage over baseline models on both CIFAR-10 and CIFAR-100. As shown in Table 1, ELA achieves 93.13% and 72.07% top-1 accuracy on the ResNet-20 backbone, surpassing ResNet-20 by 1.78% and 4.12% on CIFAR-10 and CIFAR-100, respectively. ELA also outperforms several channel and spatial attention models, such as SE [13], BAM [24], and CBAM [37]. In the ResNet-56 backbone, ELA surpasses SE, BAM, and CBAM by 1.51%, 1.63%, and 1.31% on CIFAR-10, and by 1.70%, 1.84%, and 1.77% on CIFAR-100, respectively, with lower parameters introduced. Furthermore, in ResNet-110, our model exceeds a number of layer interaction models, surpassing DIA[16] and MRLA-L [8] by 0.95% and 0.94% on CIFAR-10, and by 1.11% and 1.23% on CIFAR-100, respectively. ELA also outperforms recent attention mechanisms such as CAA [2] and SHSA [39], surpassing them by 0.81 and 0.86 on CIFAR-10, and by 0.86% and 0.43% on CIFAR-100 in ResNet-20, respectively. In ResNet-56 and ResNet-110, ELA also exceeds Dynamic Layer Attention (DLA) [33], achieving improvements of 0.37% and 0.72% in CIFAR-10 and 0.69% and 0.76% in CIFAR-100. These results highlight ELA’s superior classification performance while maintaining efficient parameter usage across various network depths.
@@
ELA significantly reduces training time. As shown in Figure 4, the training time for the five layer attention networks on CIFAR-10 and CIFAR-100 are nearly identical for one epoch, with DLA exhibiting the longest training time. Compared to MRLA-B, MRLA-L reduces training time by approximately 10%-15%, while the proposed ELA achieves an even greater reduction, decreasing training time by around 30%-35% relative to the original MRLA-B. ELA also reduces training time by 5% to 7% compared to DIA, the most lightweight layer interaction network to date. Moreover, the reduction in training time of ELA becomes more significant as the number of network layers increases.
@@
| DLA | 47.8M | 8.1B | 78.8 | 94.5 |
| ELA | 44.9M | 8.1B | 79.1 | 94.6 |
@@
Results on ImageNet-1K We compare ELA with a range of state-of-the-art models, using ResNet as backbone. As shown in Table 2, ELA achieves 78.1% Top-1 accuracy on the ResNet-50, an improvement of 2.0% over its backbone. Compared to attention-based models like SE and CBAM, ELA offers 1.4% and 0.8% higher Top-1 accuracy, respectively, while using a smaller number of parameters. Moreover, ELA outperforms [5] by 1.1%, AA [1] by 0.4%, and GC [3] by 0.4%. Furthermore, ELA performs better than MRLA-L, which achieves 77.7%, by 0.4% in the accuracy of Top-1. For the deeper ResNet-101 backbone, ELA again delivers the best performance, achieving 79.1% in Top-1 accuracy, surpassing ResNet-101 by 1.7%. It also outperforms SE, CBAM, and AA by 0.5%, 0.6%, and 0.4%, respectively. Notably, ELA achieves 0.4% higher Top-1 accuracy than MRLA-L, which reaches 78.7%. When compared to lightweight models such as ECA, ELA provides an additional 0.6% improvement in Top-1 accuracy for ResNet-50 and 0.4% for ResNet-101, despite a slight increase in parameters. Moreover, compared to the latest layer interaction networks DLA and S6LA [20], ELA achieves performance gains of 0.3% and 0.1% on ResNet-50, respectively. The results demonstrate that ELA significantly outperforms other models in terms of Top-1 and Top-5 accuracy, while maintaining competitive computational efficiency.
4.2 Object Detection
Experimental Settings In the object detection task, our approach was evaluated on the COCO2017 dataset using Faster R-CNN and Mask R-CNN as detectors. All models were implemented using the open-source MMDetection toolkit [4] to ensure consistency and reproducibility, with all configurations following the default MMDetection settings. Each detector underwent 12 training epochs, with KL divergence measured at the 1st epoch and 7th epoch to analyze layer significance at early and later training stages. The hyperparameters and were set to 2 and 5, respectively, while was fixed at 0.2.
Results on COCO2017 Our proposed ELA method achieves significant improvements in object detection performance compared to several state-of-the-art methods across various backbone networks and detectors. As highlighted in Table 3, when Faster R-CNN is used with ResNet-50, ELA outperforms SE, ECA, and RLA [41] by achieving 40.7% in AP, a gain of 3.0% over SE and 2.7% over ECA and 1.9% over RLA. It also shows improvements in AP50 and AP75, highlighting its superiority in detecting objects across different scales. With ResNet-101, ELA continues to show competitive performance, surpassing not only MRLA-L, but also SE, ECA, and other models, with a 0.1% improvement in AP and a notable increase in AP50 and AP75. Similarly, when Mask R-CNN is used as the detector, ELA maintains its edge, achieving a 1.0% increase in AP compared to MRLA-L on ResNet-50, and showing substantial improvements over methods like NL [35] and GC [3]. Moreover, ELA achieves state-of-the-art performance compared to recent layer interaction networks, such as DLA and S6LA, on both Faster R-CNN and Mask R-CNN, while maintaining the lowest parameter count. These results confirm that ELA consistently enhances object detection accuracy, offering a balanced trade-off between performance and computational efficiency.
@@ 5.1 Evaluation of Different Layer Pruning Methods
To evaluate the effectiveness of EBQM, we conducted comparative experiments with two alternative strategies. First, a fixed layer pruning strategy was employed, where the number of pruned layers was determined by a hyperparameter. The pruned layer count was aligned as closely as possible with the number determined by Algorithm 1. Second, a threshold-based pruning strategy was implemented, using a fixed hyperparameter to define a KL divergence threshold: layers with values below the threshold were discarded, while those above were retained. The ablation experiments were conducted on CIFAR-100 using ResNet [10] architectures of varying depths as baselines. Each experiment was repeated five times, and results were reported as Accuracy Standard Deviation to evaluate the model’s robustness and performance.
[…]
| Fixed Layer Pruning | 74.290.30 | |
| EBQM | 75.480.24 | |
[…]
| Fixed Layer Pruning | 74.920.41 | |
[…]
As shown in Table 4, it is evident that as the network depth increases, the instability of the Threshold-Based method also increases. The standard deviation rises from 0.44% in ResNet-56 to 0.79% in ResNet-110. A similar trend is observed with the Fixed Layer Pruning method. In contrast, the Top-1 accuracy achieved by our EBQM is consistently higher than that of the other two methods, reaching 75.48% and 76.25% for ResNet-56 and ResNet-110, respectively.
5.2 Evaluation of Different Quantile Mapping Algorithms
To evaluate the performance of EBQM in comparison to existing quantile mapping algorithms, we compared EBQM with widely used gamma quantile mapping (GQM) methods [25] and exponential quantile mapping (EQM) methods [12].
[…]
| EQM | 73.930.34 | |
| EBQM | 75.480.24 | |
[…]
| EQM | 75.050.37 | |
[…]
EBQM surpasses both GQM and EQM in Top-1 accuracy, achieving 75.48% and 76.25% for ResNet-56 and ResNet-110, respectively, as demonstrated in Table 5. These results surpass those of GQM, with accuracies of 74.02% and 74.98%, and EQM, with accuracies of 73.93% and 75.05%. Moreover, EBQM demonstrates the lowest standard deviation, at 0.24% and 0.22%, indicating superior stability when compared to GQM and EQM.
@@ 5.3 Evaluation of Different Distributions and Activation Functions in EBQM
The Beta distribution outperforms both the Normal and Gamma distributions, as well as the Softmax and Sigmoid functions, in terms of Top-1 accuracy, achieving 75.48% and 76.25% for ResNet-56 and ResNet-110, respectively, as shown in Table 6. Furthermore, the Beta distribution exhibits superior stability, with the lowest standard deviations of 0.24% and 0.22%, indicating more consistent performance compared to the other distributions.
@@ Appendix A Quantile Mapping
Algorithm 2 outlines the original Quantile Mapping Algorithm as introduced by [26]. In this method, the Gamma distribution and Exponential distribution are frequently utilized to model the target distribution . Their cumulative distribution functions (CDFs) are as follows: